what is the volume of a cylinder with a radius of 6cm, and a height of 3cm?
Answer by checkley77(12569)

AREA OF BASE=PIR^2
AREA=3.14*6^2
AREA=3.14*36
AREA=113.04 cm^2 ANSWER FOR THE BASE.
113.04*3=339.12 IN^3 IS THE VOLUME OF THE CYLINDER.